# Are Language Models Actually Useful for Time Series Forecasting?

### v1

@@ E.1 Confidence Intervals for Forecasting
Since LLMs and deep learning models in general are probabilistic in nature, their predictions can vary across different runs and different random initializations. Thus, we report the confidence intervals (CIs) for all MAE and MSE predictions made by our baseline models. We report the CIs for MAE prediction by Time-LLM, LLaTA, OneFitsAll in Tables 911, and 13, for MSE predictions in Tables 1012, and 14, respectively. Across all results, we note that the range of variation is quite small, and these intervals do not affect our observations. To illustrate the subtle differences more clearly, we present the visualized results in Figure 5 and Figure 6.
@@ E.2 Complete Results
Randomized Parameters and Encoder Exploration. The results of the randomized parameters are shown in Table 18. The results of the encoders’ exploration are shown in Table 19.
Random-Shuffling of Inputs. The remaining results for shuffled input are shown in Table 20.
@@
Models LLaTA w/o LLM LLM2Attn LLM2Trsf From Paper
@@
Table 11: Confidence Intervals for MAE predictions of LLaTA. The best performing model in highlighted in Red color text. #Wins refers to the total number of times the method performed best.
Models LLaTA w/o LLM LLM2Attn LLM2Trsf From Paper
@@
Table 12: Confidence Intervals for MSE predictions of LLaTA. The best performing model in highlighted in Red color text. #Wins refers to the total number of times the method performed best. Note that ’-’ means the dataset has not been included in the original paper.
@@
Models LLaTA w/o LLM LLM2Attn LLM2Trsf
@@
Models LLaTA w/o LLM LLM2Attn LLM2Trsf
@@
Table 16: Confidence Intervals for MAE and MSE predictions of LLaTA. The best performing model in highlighted in Red color text. #Wins refers to the total number of times the method performed best. The ablation results in the table above are from datasets that have not been studied by the reference methods [42, 13, 19].
@@
Methods LLaTA [19] OneFitAll [42] Time-LLM [13] D-LTrsf LTrsf PAttn D-PAttn DLinear MeanP Seasonal
$1^{st}$ and $2^{nd}$ Wins 33 3 2 34 21 3 0 0
$1^{st}$ and $2^{nd}$ Wins 11 0 16 4 1 0 0 0
@@
LLaTA [19] Dataset ETTh1 Illness
@@
Table 20: Results for input shuffling/masking for Time-LLM, LLaTA, and OneFitsAll methods on ETTh1 (predict length are "96, 192, 336 and 720") and Illness (predict length are "24, 36, 48 and 60"), the impact of shuffling the input on the degradation of time series forecasting performance does not change significantly before and after model modifications.
Figure 7: Ablation methods consume less time for inference while providing better forecasting performance in most cases. The figure above shows the inference time and prediction accuracy of Time-LLM, OneFitsAll, and LLaTA on ETTh1, ETTh2, ETTm1, Illness, and Weather, Traffic datasets, averaged across prediction lengths. Results of other datasets refer to Figure 3.
@@
Figure 8: Ablation methods consume less time for inference while providing better forecasting performance in most cases. The figure above shows the inference time and prediction accuracy of Time-LLM, OneFitsAll, and LLaTA on all the datasets, averaged across prediction lengths in MSE metric.

### v2

@@ E.1 Confidence Intervals for Forecasting
Since LLMs and deep learning models in general are probabilistic in nature, their predictions can vary across different runs and different random initializations. Thus, we report the confidence intervals (CIs) for all MAE and MSE predictions made by our baseline models. We report the CIs for MAE prediction by Time-LLM, CALF, OneFitsAll in Tables 911, and 13, for MSE predictions in Tables 1012, and 14, respectively. Across all results, we note that the range of variation is quite small, and these intervals do not affect our observations. To illustrate the subtle differences more clearly, we present the visualized results in Figure 5 and Figure 6.
@@ E.2 Complete Results
Randomized Parameters and Random-Shuffling of Inputs. The results of the randomized parameters are shown in Table 20. The remaining results for shuffled input are shown in Table 21,22,23 and 24.
@@
Models CALF w/o LLM LLM2Attn LLM2Trsf From Paper
@@
Table 11: Confidence Intervals for MAE predictions of CALF. The best performing model in highlighted in Red color text. #Wins refers to the total number of times the method performed best.
Models CALF w/o LLM LLM2Attn LLM2Trsf From Paper
@@
Table 12: Confidence Intervals for MSE predictions of CALF. The best performing model in highlighted in Red color text. #Wins refers to the total number of times the method performed best. Note that ’-’ means the dataset has not been included in the original paper.
@@
Models CALF w/o LLM LLM2Attn LLM2Trsf
@@
Models CALF w/o LLM LLM2Attn LLM2Trsf
@@
Table 16: Confidence Intervals for MAE and MSE predictions of CALF. The best performing model in highlighted in Red color text. #Wins refers to the total number of times the method performed best. The ablation results in the table above are from datasets that have not been studied by the reference methods [49, 14, 21].
@@
Methods CALF [21] OneFitAll [49] Time-LLM [14] D-LTrsf LTrsf PAttn D-PAttn DLinear MeanP Seasonal
$1^{st}$ and $2^{nd}$ Wins 33 3 2 34 21 3 0 0
$1^{st}$ and $2^{nd}$ Wins 11 0 16 4 1 0 0 0
Simple Methods LLM-Based Methods Other State-of-the-art Methods
Methods LTrsf PAttn CALF [21] OneFitsAll [49] Time-LLM [14] DLinear [45] RLinear [20] PatchTST [26] iTransformer [23] TimesNet [40] FEDformer [48]
Metric MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
ETTm1 96 0.351 0.300 0.342 0.301 0.349 0.323 0.346 0.292 0.334 0.272 0.343 0.299 0.376 0.355 0.342 0.290 0.368 0.334 0.375 0.338 0.419 0.379
192 0.366 0.327 0.366 0.344 0.375 0.374 0.372 0.332 0.358 0.310 0.365 0.335 0.392 0.391 0.369 0.332 0.391 0.377 0.387 0.374 0.441 0.426
336 0.386 0.356 0.383 0.369 0.399 0.409 0.394 0.366 0.384 0.352 0.386 0.369 0.415 0.424 0.392 0.366 0.420 0.426 0.411 0.410 0.459 0.445
720 0.416 0.421 0.413 0.426 0.438 0.477 0.421 0.417 0.411 0.383 0.421 0.425 0.450 0.487 0.420 0.416 0.459 0.491 0.450 0.478 0.490 0.543
ETTm2 96 0.258 0.172 0.249 0.164 0.256 0.178 0.262 0.173 0.253 0.161 0.260 0.167 0.265 0.182 0.255 0.165 0.264 0.180 0.267 0.187 0.287 0.203
192 0.295 0.226 0.292 0.224 0.297 0.242 0.301 0.229 0.293 0.219 0.303 0.224 0.304 0.246 0.292 0.220 0.309 0.250 0.309 0.249 0.328 0.269
336 0.323 0.269 0.319 0.266 0.339 0.307 0.341 0.286 0.392 0.271 0.343 0.281 0.342 0.307 0.329 0.274 0.348 0.311 0.351 0.321 0.366 0.325
720 0.372 0.344 0.370 0.345 0.393 0.397 0.401 0.378 0.379 0.352 0.421 0.397 0.398 0.407 0.385 0.362 0.407 0.412 0.403 0.408 0.415 0.421
ETTh1 96 0.390 0.371 0.387 0.383 0.389 0.369 0.397 0.376 0.392 0.362 0.399 0.375 0.395 0.386 0.399 0.370 0.405 0.386 0.402 0.384 0.419 0.376
192 0.418 0.428 0.416 0.439 0.423 0.427 0.418 0.416 0.418 0.398 0.416 0.405 0.424 0.437 0.421 0.413 0.436 0.441 0.429 0.436 0.448 0.420
336 0.426 0.449 0.419 0.447 0.436 0.456 0.433 0.442 0.427 0.430 0.443 0.439 0.446 0.479 0.436 0.422 0.458 0.487 0.469 0.491 0.465 0.459
720 0.465 0.485 0.457 0.474 0.467 0.479 0.456 0.477 0.457 0.442 0.490 0.472 0.470 0.481 0.466 0.447 0.491 0.503 0.500 0.521 0.507 0.506
ETTh2 96 0.333 0.285 0.327 0.277 0.331 0.279 0.342 0.285 0.328 0.268 0.353 0.289 0.338 0.288 0.336 0.274 0.349 0.297 0.374 0.340 0.397 0.358
192 0.376 0.351 0.373 0.345 0.380 0.353 0.389 0.354 0.375 0.329 0.418 0.383 0.390 0.374 0.379 0.339 0.400 0.380 0.414 0.402 0.439 0.429
336 0.379 0.346 0.375 0.338 0.394 0.362 0.407 0.373 0.409 0.368 0.465 0.448 0.426 0.415 0.380 0.329 0.432 0.428 0.452 0.452 0.487 0.496
720 0.426 0.406 0.425 0.403 0.426 0.404 0.441 0.406 0.420 0.372 0.551 0.605 0.440 0.420 0.422 0.379 0.445 0.427 0.468 0.462 0.474 0.463
Electricity 96 0.220 0.130 0.226 0.134 0.238 0.145 0.238 0.139 0.224 0.131 0.237 0.140 0.281 0.201 0.222 0.129 0.240 0.148 0.272 0.168 0.308 0.193
192 0.239 0.149 0.239 0.149 0.252 0.161 0.251 0.153 0.241 0.152 0.249 0.153 0.283 0.201 0.240 0.157 0.253 0.162 0.289 0.184 0.315 0.201
336 0.254 0.164 0.257 0.166 0.267 0.175 0.266 0.169 0.248 0.160 0.267 0.169 0.298 0.215 0.259 0.163 0.269 0.178 0.300 0.198 0.329 0.214
720 0.278 0.193 0.288 0.204 0.303 0.222 0.297 0.206 0.298 0.192 0.301 0.203 0.331 0.257 0.290 0.197 0.317 0.225 0.320 0.220 0.355 0.246
Traffic 96 0.234 0.357 0.262 0.394 0.268 0.407 0.282 0.388 0.248 0.362 0.282 0.410 0.389 0.649 0.249 0.360 0.268 0.395 0.321 0.593 0.366 0.587
192 0.243 0.377 0.265 0.404 0.278 0.430 0.290 0.407 0.247 0.374 0.287 0.423 0.366 0.601 0.256 0.379 0.276 0.417 0.336 0.617 0.373 0.604
336 0.250 0.388 0.270 0.413 0.281 0.444 0.294 0.412 0.271 0.385 0.296 0.436 0.369 0.609 0.264 0.392 0.283 0.433 0.336 0.629 0.383 0.621
720 0.273 0.428 0.285 0.448 0.300 0.477 0.312 0.450 0.288 0.430 0.315 0.466 0.387 0.647 0.286 0.432 0.302 0.467 0.350 0.640 0.382 0.626
Weather 96 0.204 0.161 0.191 0.151 0.204 0.164 0.212 0.162 0.201 0.147 0.237 0.176 0.232 0.192 0.198 0.149 0.214 0.174 0.220 0.172 0.296 0.217
192 0.246 0.205 0.232 0.193 0.250 0.214 0.248 0.204 0.234 0.189 0.282 0.220 0.271 0.240 0.241 0.194 0.254 0.221 0.261 0.219 0.336 0.276
336 0.283 0.253 0.273 0.244 0.291 0.269 0.286 0.254 0.279 0.262 0.319 0.265 0.307 0.292 0.282 0.245 0.296 0.278 0.306 0.280 0.380 0.339
720 0.332 0.322 0.323 0.311 0.352 0.355 0.337 0.326 0.316 0.304 0.362 0.323 0.353 0.364 0.334 0.314 0.347 0.358 0.359 0.365 0.428 0.403
$1^{st}$ Wins 11 18 0 1 25 0 0 4 0 0 0
Table 19: Comparison between simple ablations, LLM-based methods, and non-LLM methods. LLM-based methods slightly outperform non-LLM methods. The non-LLM results are from the LLM-Time and iTransformer papers, which are computed on the same splits. As expected, the main findings in our paper are unchanged: The LLM-based methods are slightly better, but our ablations indicate this isn’t due to the LLM. Red: Best performance. Blue: Second Best.
@@
CALF [21] Dataset ETTh1 Illness
@@
Table 21: Results for input shuffling/masking for Time-LLM, CALF, and OneFitsAll methods on ETTh1 (predict length are "96, 192, 336 and 720") and Illness (predict length are "24, 36, 48 and 60"), the impact of shuffling the input on the degradation of time series forecasting performance does not change significantly before and after model modifications.
Dataset ETTh2 Electricity
Time-LLM 27.1% 4.5% 44.6% 99.5% 212.1% 52.2% 323.4% 103.3%
w/o LLM 31.2% 2.9% 52.4% 137.9% 220.9% 47.3% 332.4% 105.5%
LLM2Attn 30.7% 2.9% 46.7% 134.8% 234.7% 58.2% 350.4% 115.3%
LLM2Trsf 24.8% 2.3% 43.4% 115.0% 240.8% 50.8% 362.0% 118.5%
OneFitsAll 33.0% 1.1% 39.7% 112.1% 237.8% 25.2% 382.0% 88.9%
w/o LLM 32.7% 2.8% 38.7% 115.8% 249.7% 43.4% 384.7% 109.2%
LLM2Attn 27.5% 1.0% 36.5% 137.4% 255.2% 49.5% 393.9% 119.2%
LLM2Trsf 29.6% 2.2% 37.2% 130.8% 263.8% 55.1% 406.4% 125.4%
CALF 11.3% 2.0% 9.1% 44.7% 235.1% 108.2% 29.3% 37.6%
w/o LLM 23.1% 4.1% 17.2% 46.9% 239.5% 109.2% 28.6% 39.4%
LLM2Attn 16.7% 3.4% 14.8% 31.7% 237.9% 116.2% 28.9% 42.2%
LLM2Trsf 18.9% 3.0% 15.3% 43.2% 245.6% 107.0% 29.2% 42.3%
Table 22: For the input shuffling/masking experiments on ETTh2 and Electricity, the impact of shuffling the input on the degradation of time series forecasting performance does not change significantly before and after model modifications.
Dataset ETTm1 ETTm2
Time-LLM 66.6% 10.1% 107.7% 43.5% 47.0% 5.3% 77.7% 202.9%
w/o LLM 68.7% 12.4% 112.0% 52.6% 49.8% 4.7% 78.7% 199.6%
LLM2Attn 73.5% 13.5% 119.3% 50.7% 49.6% 4.7% 75.9% 196.3%
LLM2Trsf 72.0% 27.6% 117.0% 54.2% 46.4% 3.7% 76.5% 191.9%
OneFitsAll 74.4% 8.0% 123.4% 41.8% 48.0% 3.7% 82.9% 172.9%
w/o LLM 66.7% 10.3% 115.1% 45.0% 48.1% 4.4% 81.9% 190.7%
LLM2Attn 73.9% 11.1% 124.8% 51.6% 45.5% 3.3% 77.8% 183.6%
LLM2Trsf 70.1% 8.3% 126.2% 50.5% 48.2% 4.3% 82.0% 182.9%
CALF 64.3% 24.8% 122.3% 13.7% 23.5% 9.1% 47.1% 61.7%
w/o LLM 66.7% 26.0% 123.6% 15.9% 27.1% 11.2% 51.9% 58.1%
LLM2Attn 62.5% 23.4% 122.3% 15.6% 25.2% 8.8% 51.5% 57.3%
LLM2Trsf 61.8% 25.6% 122.6% 13.6% 23.7% 8.2% 51.0% 59.3%
Table 23: For the input shuffling/masking experiments on ETTm1 and ETTm2, the impact of shuffling the input on the degradation of time series forecasting performance does not change significantly before and after model modifications.
Dataset Weather Traffic
Time-LLM 65.8% 5.5% 85.9% 85.9% 198.1% 56.6% 309.8% 101.3%
w/o LLM 59.2% 4.7% 78.7% 74.6% 196.8% 70.2% 282.1% 78.3%
LLM2Attn 71.8% 8.5% 91.6% 95.3% 212.3% 82.4% 312.7% 92.2%
LLM2Trsf 71.9% 7.1% 94.9% 103.2% 197.5% 64.6% 307.2% 101.1%
OneFitsAll 71.0% 9.8% 102.1% 97.7% 206.4% 11.7% 369.1% 93.5%
w/o LLM 56.4% 2.3% 81.2% 73.5% 238.4% 67.0% 338.6% 85.9%
LLM2Attn 64.9% 5.4% 90.2% 94.8% 222.6% 57.5% 350.5% 95.1%
LLM2Trsf 77.8% 8.3% 101.9% 111.0% 217.8% 35.9% 361.3% 113.3%
CALF 32.6% 4.2% 54.5% 27.2% 201.7% 105.0% 63.1% 27.2%
w/o LLM 28.1% 5.9% 49.4% 13.6% 217.3% 115.4% 67.2% 29.7%
LLM2Attn 26.0% 6.2% 49.2% 17.1% 226.7% 122.7% 69.4% 30.7%
LLM2Trsf 28.0% 6.4% 54.1% 21.1% 232.8% 120.5% 72.0% 29.6%
Table 24: For the input shuffling/masking experiments on Weather and Traffic, the impact of shuffling the input on the degradation of time series forecasting performance does not change significantly before and after model modifications.
Figure 7: Ablation methods consume less time for inference while providing better forecasting performance in most cases. The figure above shows the inference time and prediction accuracy of Time-LLM, OneFitsAll, and CALF on ETTh1, ETTh2, ETTm1, Illness, and Weather, Traffic datasets, averaged across prediction lengths. Results of other datasets refer to Figure 3.
@@
Figure 8: Ablation methods consume less time for inference while providing better forecasting performance in most cases. The figure above shows the inference time and prediction accuracy of Time-LLM, OneFitsAll, and CALF on all the datasets, averaged across prediction lengths in MSE metric.